OpenStudy (anonymous):

-3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

-3(-3x-7y=-16) have : 9x-21y=48 Sum : -9x + 5y = 16 and 9x-21y=48 -9x+9x+5y-21y=64 have : -16y=64 y=-64/16 y=-4 x : -3x + 7y = -16 -3x-28=-16 -3x=28-16 -3x=12 x=-4 Got it ?;)

OpenStudy (anonymous):

Got it! Thank you for showing your work!! This really really helped! Thank you :)